OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57

OpenStudy (anonymous):

ummm is this algebra 2?

OpenStudy (anonymous):

algebra 1, 2nd year taking it again

OpenStudy (anonymous):

once i find the y , i put it in the the first equation right?

OpenStudy (anonymous):

I think so... i asked somebody if they could come and help since I need to understand this better as well since i sort of forgot haha :)

OpenStudy (anonymous):

sorry hold on lemme see if that person will come... ;/

OpenStudy (anonymous):

okay thank you (:

OpenStudy (anonymous):

Hi @jim_thompson5910 thanks for coming :) I also do not understand this 100% haha :) u got my message then i take it? :)

jimthompson5910 (jim_thompson5910):

double everything in the second equation to get 8x + 2y = 114

OpenStudy (anonymous):

sorry (email)

OpenStudy (anonymous):

and hi @apoorvk thanks for coming also!!! i appreciate it :)

OpenStudy (apoorvk):

ohkay, @iheartfood - go on and post what you would have done in this problem, and don't worry if it's right or wrong.. :)

jimthompson5910 (jim_thompson5910):

Now add the two equations 7x - 2y = 21 8x + 2y = 114 -------------- 15x + 0y = 135 Now solve 15x = 135 for x

OpenStudy (anonymous):

well i know you're supposed to solve for one, but i did not know you could do that (what @jim_thompson5910 wrote.... so that is another way you can do it??

OpenStudy (anonymous):

is that how you do it?? simplify it into one larger equation??

jimthompson5910 (jim_thompson5910):

that's one way to do it, there is more than one way

OpenStudy (anonymous):

okay, what's another way to do it?? i think i remember that you can solve one equation and then plug it in but i couldn't do it for this one... ;/

jimthompson5910 (jim_thompson5910):

you isolate one variable in terms of the other then you substitute 4x + y = 57 y = 57 - 4x ------------------ 7x - 2y = 21 7x - 2(57-4x) = 21 Now solve for x

OpenStudy (anonymous):

x = 9 for my problem right? (:

OpenStudy (anonymous):

oh okay cool :) soooo x=45??

jimthompson5910 (jim_thompson5910):

you are correct, the answer to the first piece is x = 9

OpenStudy (anonymous):

oops my bad nvm lol

jimthompson5910 (jim_thompson5910):

hmm how did you get x = 45?

jimthompson5910 (jim_thompson5910):

now that you know that x = 9, use this to find y

OpenStudy (anonymous):

sorry by bad... lol it's 9 sorry!!

OpenStudy (anonymous):

sooo you plug it in right??

OpenStudy (anonymous):

so y=21 ???? and x=9 ???

jimthompson5910 (jim_thompson5910):

yes, pick any equation dealing with y and plug in x = 9 7x - 2y = 21 7(9) - 2y = 21 Solve for y

jimthompson5910 (jim_thompson5910):

and yes, y = 21

OpenStudy (anonymous):

oh i used the second equation.. haha but it works both ways :) u get the same answer for both right?

jimthompson5910 (jim_thompson5910):

oh the question was just "What is the value of the x variable in the solution to the following system of equations?" so they just want x Usually, they'll want both x and y

jimthompson5910 (jim_thompson5910):

yes either equation works

OpenStudy (anonymous):

oh okay cool! thanks!! i totally understand it now... :) what a refresher!! And @sommersquid did that help??? And thanks again for coming @jim_thompson5910 :)

OpenStudy (anonymous):

yesss, it really did (: i just needed help with the second step but i got it now (: thank you so much everyone

jimthompson5910 (jim_thompson5910):

yw